if the area of a trapazoid was 72 ft^2 and the sum of the bases was 48 ft what would the heighth be?
Answer by Theo(13342)

the formula for the area of a trapezoid is 1/2 times the sum of the bases times the height.
here's a reference for that.
http://www.mathgoodies.com/lessons/vol1/area_trapezoid.html
you are given that the area is 72 square feet and you are given that the sum of the bases if 48 feet.
plug that in the formula and you get:
formula is:
area = 1/2 * height * sum of bases
this becomes:
72 = 1/2 * height * 48 which becomes:
72 = 24 * height
divide both sides of this equation by 24 to get:
72/24 = height
this makes height equal to 3 feet.
the bases could be any length as long as their sum is equal to 48.